# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ B.3 Convergence Results for AdaGrad-Norm and AdaGrad
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\frac{2L_{0}^{2}}{\sqrt{v_{k-1}}}\left\lVert x_{k}-x_{k-1}\right\rVert^{2}+\frac{2L_{1}^{2}}{\sqrt{v_{k-1}}}\left\lVert\nabla F(x_{k})\right\rVert^{2}\left\lVert x_{k}-x_{k-1}\right\rVert^{2}+\mathbb{E}_{k}[\varphi_{k-1}-\varphi_{k}].$
Again, by the A.M.-G.M. inequality, with any $\rho_{2}\in\mathbb{R}_{++}$,
@@
Then, (B.13) is further upper bounded by
@@
Appendix C Details and Additional Results of Numerical Experiments
C.1 Multi-class Logistic Regression on MNIST
@@
Optimizer scaling rule None
Base learning rate 0.008
Base batch size 2
Base weight decay 0
Weight decay scaling rule None
Momentum 0
Momentum scaling rule None
Precision tf32
C.2 Three-layer Convolutional Neural Network on MNIST
Table 3: Training hyperparameters for three-layer CNN on MNIST
Model 3-layer CNN on MNIST
Weight initialization Default
Learning rate schedule None
Training samples 6M (100 epochs)
Optimizer SGD or AdaGrad(-Norm)
Optimizer scaling rule None
@@ C.2 Three-layer Convolutional Neural Network on MNIST
Base weight decay 0
Weight decay scaling rule None
@@
Table 4: Three-layer CNN on MNIST
Scheme test steps average batch size training loss validation accuracy
AdaSGD norm; $\eta=0.10$ 256 23547 0.79 0.83
AdaSGD norm; $\eta=0.25$ 383 15627 0.48 0.88
AdAdaGrad-Norm norm; $\eta=0.10$ 226 26567 0.88 0.83
AdAdaGrad-Norm norm; $\eta=0.25$ 435 13831 0.54 0.87
AdAdaGrad norm; $\eta=0.10$ 148 40555 0.16 0.96
AdAdaGrad norm; $\eta=0.25$ 204 29480 0.12 0.97
AdaSGD inner product; $\vartheta=0.01$ 230 26074 0.98 0.80
AdaSGD inner product; $\vartheta=0.05$ 413 14593 0.45 0.88
AdAdaGrad-Norm inner product; $\vartheta=0.01$ 242 24827 0.82 0.84
AdAdaGrad-Norm inner product; $\vartheta=0.05$ 525 11458 0.50 0.88
@@
C.3 Three-layer Convolutional Neural Network on CIFAR-10
@@
Base weight decay 0
Weight decay scaling rule None
@@
Scheme test steps average batch size training loss validation accuracy
@@
Appendix D Further Related Work
D.1 Convergence of Stochastic Gradient Methods
We give a brief overview on the convergence results of SGD and adaptive gradient methods. The convergence of SGD for smooth nonconvex functions was first analyzed in Ghadimi and Lan (2013), assuming a uniform bound on the variance of stochastic gradients. Arjevani et al. (2023) gave a tight lower bound. Bottou et al. (2018) extended the convergence result with the so-called affine variance noise model. Different assumptions on the second moment of stochastic gradients appearing in the literature were reviewed in Khaled and Richtárik (2023), with a view to developing a general convergence theory of SGD for smooth nonconvex functions.

### v2

@@ C.3 Convergence Results for AdaGrad-Norm and AdaGrad
$\qquad\qquad+\frac{2L_{0}^{2}}{\sqrt{v_{k-1}}}\left\lVert x_{k}-x_{k-1}\right\rVert^{2}+\frac{2L_{1}^{2}}{\sqrt{v_{k-1}}}\left\lVert\nabla F(x_{k})\right\rVert^{2}\left\lVert x_{k}-x_{k-1}\right\rVert^{2}+\mathbb{E}_{k}[\varphi_{k-1}-\varphi_{k}].$
Again, by the A.M.-G.M. inequality, with any $\rho_{2}>0$,
@@
Then, (C.14) is further upper bounded by
@@
Appendix D Details and Additional Results of Numerical Experiments
For the implementation of the proposed schemes, per-sample gradients are computed using JAX-like composable function transforms called torch.func in PyTorch 2.0+. Numerical experiments are carried out on workstations with NVIDIA RTX 2080Ti 11GB (for MNIST), A100 80GB (for CNN on CIFAR-10) and L40S 48GB (for ResNet-18 on CIFAR-10) GPUs, with PyTorch 2.2.1 (Paszke et al., 2019) and Lightning Fabric 2.2.0 (Falcon and The PyTorch Lightning team, 2019). The AdaGrad-Norm implementation is taken from (Ward et al., 2019).
D.1 Multi-class Logistic Regression on MNIST
We initially applied our methods to a ten-class logistic regression problem using the MNIST dataset, with a smooth convex objective. Our experiments were conducted with an equal training budget of 6 million samples (equivalent to 100 epochs), setting a maximum batch size of 60,000 (i.e., the full batch) for all approaches. To highlight the adaptivity and flexibility of our proposed methods, we refrained from conducting an exhaustive search for optimal values of $\alpha$, $\eta$, and $\vartheta$, and we did not employ any learning rate schedules across the methods. The outcomes, including the number of iterations required (steps), average batch sizes (bsz.), final training loss (loss), and final validation accuracy (acc.), are documented in Table 3.
Table 3: Multi-class logistic regression on MNIST
Scheme test steps bsz. loss acc.
AdaSGD $\eta=0.10$ 351 17131 1.04 0.82
AdaSGD $\eta=0.25$ 1029 5831 0.67 0.86
AdAdaGrad-Norm $\eta=0.10$ 596 10060 1.50 0.78
AdAdaGrad-Norm $\eta=0.25$ 2462 2437 1.02 0.83
AdAdaGrad $\eta=0.10$ 126 47282 0.54 0.88
AdAdaGrad $\eta=0.25$ 274 21918 0.46 0.90
AdaSGD $\vartheta=0.05$ 717 8362 0.76 0.85
AdaSGD $\vartheta=0.10$ 1804 3327 0.54 0.88
AdAdaGrad-Norm $\vartheta=0.05$ 1127 5322 1.28 0.80
AdAdaGrad-Norm $\vartheta=0.10$ 5349 1122 0.80 0.85
The results, coupled with the graphical representations in Figure 4, reveal that AdAdaGrad, utilizing the norm test with $\eta=0.25$, outperforms others in terms of both training loss and validation accuracy. These findings underscore the importance of adaptive (and coordinate-wise) learning rates within adaptive sampling methods. Despite the best performance of AdAdaGrad using the norm test with $\eta=0.25$, we note that the one with $\eta=0.10$ is more hardware-efficient—having an average batch size of more than 47,000, totaling only 126 steps (i.e., gradient evaluation) to run through all 6 million samples. This implies an intrinsic trade-off between computationally efficiency and generalization. Interestingly, AdAdaGrad-Norm, while theoretically simpler to analyze, might underperform compared to AdaSGD (i.e., SGD with adaptive batch size schemes) for this specific convex problem and set of hyperparameters. It is also noteworthy that the norm test tends to increase batch sizes more aggressively than the inner product test, leading to more efficient utilization of available GPU memory.
@@ D.1 Multi-class Logistic Regression on MNIST
Optimizer SGD or AdaGrad(-Norm)
Base learning rate 0.008
Base batch size 2
Maximum global batch size 60000
Weight decay 0
Momentum 0
Precision tf32
D.2 Three-layer Convolutional Neural Network on MNIST
Table 5: Three-layer CNN on MNIST
Scheme test steps time (h) bsz. loss acc.
SGD N/A 2929 6.77 2048 0.12 0.96
SGD N/A 1464 3.38 4096 0.20 0.94
SGD N/A 732 1.72 8192 0.32 0.91
SGD N/A 366 0.90 16384 0.51 0.87
SGD N/A 183 0.45 32768 1.54 0.75
SGD N/A 99 0.27 60000 2.15 0.66
AdaGrad N/A 2929 7.12 2048 0.02 0.99
AdaGrad N/A 1464 3.60 4096 0.02 0.99
AdaGrad N/A 732 1.82 8192 0.05 0.98
AdaGrad N/A 366 0.92 16384 0.07 0.98
AdaGrad N/A 199 0.52 30000 0.10 0.97
AdaGrad N/A 183 0.47 32768 0.11 0.97
AdaGrad N/A 149 0.36 40000 0.13 0.96
AdaGrad N/A 99 0.29 60000 0.17 0.95
AdaSGD norm; $\eta=0.10$ 256 0.73 23546 0.79 0.83
AdaSGD norm; $\eta=0.25$ 383 1.05 15627 0.48 0.88
AdAdaGrad-Norm norm; $\eta=0.10$ 226 0.65 26567 0.88 0.83
AdAdaGrad-Norm norm; $\eta=0.25$ 435 1.27 13830 0.54 0.87
AdAdaGrad norm; $\eta=0.10$ 149 0.45 40057 0.15 0.96
AdAdaGrad norm; $\eta=0.25$ 198 0.58 30152 0.13 0.97
AdAdaGrad norm; $\eta=0.5$ 215 0.62 27940 0.11 0.97
AdAdaGrad norm; $\eta=0.75$ 271 0.79 22228 0.10 0.97
AdaSGD inner product; $\vartheta=0.01$ 230 0.63 26078 0.98 0.80
AdaSGD inner product; $\vartheta=0.05$ 411 1.17 14593 0.45 0.88
AdAdaGrad-Norm inner product; $\vartheta=0.01$ 241 0.70 24872 0.83 0.84
AdAdaGrad-Norm inner product; $\vartheta=0.05$ 528 1.44 11365 0.50 0.88
Table 6: Training hyperparameters for three-layer CNN on MNIST
Model 3-layer CNN on MNIST
Training samples 6M (100 epochs)
Weight initialization Default
@@ D.2 Three-layer Convolutional Neural Network on MNIST
Maximum batch size 60000
Weight decay 0
@@
Figure 6: Training loss, validation accuracy and batch size curves (vs. number of training samples) of SGD and AdaSGD for three-layer CNN on the MNIST dataset.
Figure 7: Training loss, validation accuracy and batch size curves (vs. number of training samples) of AdaGrad and AdAdaGrad for three-layer CNN on the MNIST dataset.
D.3 Three-layer Convolutional Neural Network on CIFAR-10
@@
Maximum batch size 10000
Weight decay 0
@@
Scheme test steps bsz. loss acc.
@@
D.4 ResNet-18 on CIFAR-10
Table 9: Training hyperparameters for ResNet-18 on CIFAR-10
Model ResNet-18 on CIFAR-10
Training samples 10M (200 epochs)
Weight initialization Default
Optimizer AdaGrad or Adam
Learning rate schedule Linear warmup + cosine decay
Learning rate warmup (samples) 1M
$(\beta_{1},\beta_{2})$ $(0.9,0.95)$
$\varepsilon$ $10^{-8}$
Peak learning rate 0.05
Minimum learning rate 0.005
Base batch size 8
Maximum batch size 50000
Weight decay 0
Precision tf32
Table 10: ResNet-18 on CIFAR-10
Scheme test steps time (h) bsz. loss acc.
AdaGrad N/A 2441 0.88 4096 0.0042 0.8521
AdaGrad N/A 1220 0.70 8192 0.0808 0.8072
AdaGrad N/A 610 0.56 16384 0.5098 0.7264
AdaGrad N/A 305 0.32 32768 0.9684 0.5816
AdaGrad N/A 199 0.23 50000 1.3625 0.4708
Adam N/A 2441 1.20 4096 0.0003 0.9147
Adam N/A 1220 0.97 8192 0.0004 0.8946
Adam N/A 610 0.77 16384 0.0028 0.8628
Adam N/A 305 0.45 32768 0.4000 0.7463
Adam N/A 199 0.33 50000 1.0680 0.5750
AdAdaGrad norm; $\eta=0.025$ 222 0.32 44934 1.2770 0.5107
AdAdaGrad norm; $\eta=0.05$ 485 0.60 20615 0.6204 0.7079
AdAdaGrad norm; $\eta=0.075$ 1697 1.02 5892 0.0258 0.8180
AdAdaGrad norm; $\eta=0.1$ 1404 0.94 7123 0.0668 0.8085
AdAdam norm; $\eta=0.025$ 211 0.34 47380 0.9039 0.6234
AdAdam norm; $\eta=0.05$ 426 0.60 23463 0.0061 0.8228
AdAdam norm; $\eta=0.075$ 900 0.74 11108 0.0008 0.8983
AdAdam norm; $\eta=0.1$ 1126 0.82 8880 0.0000 0.9042